OpenStudy (anonymous):

inverse function of f(n)=2(n-2)^3

OpenStudy (hitaro9):

While maybe not the most conceptually easy to understand, I think the easiest way to get an answer with one of these problems is to switch the f(x) and the x So f(n) = 2(n-3)^3 may be written as n = 2(f(n)-2)^3 And then you use algebra 1 stuff to rearrange in terms of f(n).

OpenStudy (hitaro9):

You able to get the answer from that, or do you still need help?

OpenStudy (hitaro9):

If inverses aren't making a ton of sense, try checking out this series of videos. If you've never used khan academy before, it's a wonderful resource. https://www.khanacademy.org/math/algebra/algebra-functions/function_inverses/v/introduction-to-function-inverses